OpenStudy (anonymous):

solve by substitution method x+y=-4 y=-3x+2

OpenStudy (anonymous):

substituting value from eqn2 to eqn 1 x+(-3x+2) = -4 x-3x+2=-4 -2x=-6 x=3 y=-7

OpenStudy (anonymous):

x + y = -4 x = -4 - y y = -3x + 2 = -3(-4-y) + 2 = 12 + 3y + 2 -2y = 14 y = -7 x = -4 -(-7) = -4 + 7 = 3

OpenStudy (anonymous):

x=3 y=-7

OpenStudy (anonymous):

thank ya